### v3

@@ 4.3 Spatial distribution of perception
Figure 6: Collected walkability ratings in and around Amsterdam, averaged into hexagonal bins
@@ 5.1 Main findings
Our method of scoring is easier to use than that of Place Pulse (Dubey et al., 2016; Salesses, 2012) or other relative comparison studies (Naik et al., 2014; Ye et al., 2019; Larkin et al., 2022). Like Kruse et al. (2021), we collect ratings along an absolute five-point scale, whereas much of the challenge in projects such as Place Pulse was in reconstructing a ranking from a large set of relative comparisons; this is a basic philosophical difference in approach. The advantage of using absolute scores is that the ratings are always linearly ordered. We also argue that with high-level and high-complexity concepts such as ‘walkability’ and ‘safety’ the use of relative comparisons is not necessarily helpful, since there are so many aspects of the imagery to consider simultaneously; past psychological research to justify the use of relative comparisons (Stewart et al., 2005) considered only low-complexity concepts such as ‘line lengths’ or ‘sound tones’. Instead of having a complex process of translating relative comparisons into absolute scores, we keep things simple in our system by showing only one image and enabling participants to give a quick first-impression one-click/swipe answer from a selection of five possible responses. This user interface helps participants quickly and smoothly look through many more images than interfaces with a rating slider (Twedt et al., 2016; Larkin et al., 2022) and avoids any biasing problems created by feeding back preliminary machine learning model results into the user interface (Yao et al., 2019), while still allowing participants to able to enjoy rapid progression through the survey.
@@
We were able to easily attract public participants and give them the flexibility to respond at their convenience because our survey is a smartphone-friendly web app rather than a downloaded software application. We did not pay participants (Twedt et al., 2016; Pearson et al., 2024; Larkin et al., 2022; Kruse et al., 2021) but rather relied on the attractiveness of the web app as a fun ‘game’ that can be played during downtime (e.g. waiting for the bus).
In the future, we will expand our toolkit by building a framework for generating models of the data collected by our street perception surveys. The state-of-the-art in image classification has moved on considerably in the past five years. Within the past year alone, researchers have published large pre-trained models such as StreetCLIP (Haas et al., 2023), which is designed to perform zero-shot image classification and has been specialised by training on a geo-tagged SVI dataset containing 1.1 million images. For example, as input to StreetCLIP one could provide a photograph taken from a city street and a set of three descriptive captions such as ‘An urban street’, ‘A suburban road’ and ‘A rural road’. The output of the model would be a probability distribution indicating which one of the input descriptions is the most likely to be suitable for the given image. StreetCLIP is based on the CLIP text and image association models (Radford et al., 2021). We can take StreetCLIP even further by fine-tuning it on our collected images and ratings by transforming our categories and five-point numerical rating scale into textual image captions using a template such as ‘An image of a street with a [rating text] [category] rating’. For example a walkability rating of 5 becomes ‘An image of a street with a very good walkability rating’, and a safety rating of 2 becomes ‘An image of a street with a bad safety rating’. The resulting image and text models will be able to take an input of an image and a selection of five templated sentences, and then output a probability distribution indicating which of those the most likely sentence to apply to the given image. In this manner, we can extrapolate the perception data by starting from the collected ratings for tens of thousands of images and then obtain modelled ratings for millions of images. One disadvantage of StreetCLIP is that the published model is based on Google SVI, therefore one option we must consider is retraining it based on open or VSVI instead, in addition to our collected dataset.
@@ 5.2 Limitations
Mapillary777mapillary.com/about, as of February 2024 has more than 2 billion images from 190 countries, and especially good coverage in countries like the Netherlands, including almost full panoramic coverage of Amsterdam-area streets. However, although Mapillary has imagery from every continent, the most abundant SVI coverage comes from Europe and North America (Ma et al., 2019). However, commercial SVI is no panacea either: Kim and Jang (2023) analyzed Google SVI coverage of walk commute trajectories in small- to medium-sized cities in the United States and found gaps in nearly half of the routes they investigated; they suggest that researchers consider Mapillary to bolster coverage.
@@ 6 Conclusion
We offer to the research community an open-source toolkit for downloading, processing and filtering VSVI from a given geographic region, and deploying a mobile-friendly human perception survey web app on the resulting images. In the spirit of citizen science and the FAIR research principles, anyone may easily replicate, modify and deploy this toolkit on any location of interest. We integrated it with the Mapillary platform because they provide open access to VSVI from many cities around the world. This toolkit will be used to build deep learning models with the collected data and make predictions about human perception of large amounts of SVI over a wider area. Ultimately, we hope to see such perception data used in research that can help guide future planning and development choices, and therefore improve the quality of life for many people in urban areas.
@@ Appendix C Processing and filtering VSVI details
For full up-to-date usage information please see the Spatial Data Science and Geo AI Lab888github.com/Spatial-Data-Science-and-GEO-AI-Lab/percept web site.
@@ C.1 Downloading imagery from Mapillary
This script takes a Mapillary API key999www.mapillary.com/developer and a bounding box (west, south, east, north) and conducts a lengthy but robust and restartable procedure to methodologically find and download each tile data file, cache it, and then download all of the eligible SVI that is found within the tile (and within the bounding box).

### v4

@@ 4.3 Spatial distribution of perception
Figure 6: Collected walkability ratings in and around Amsterdam. The inset map shows individual points colored by score within a small region of the city, the larger map shows these points aggregated and averaged into hexagonal bins.
@@ 5.1 Main findings
Our survey app asks the participant to view only a single image at a time, unlike Place Pulse (Dubey et al., 2016; Salesses, 2012) or other relative comparison studies (Naik et al., 2014; Ye et al., 2019; Larkin et al., 2022), which involve showing the participant at least two images at a time. Like Kruse et al. (2021), we collect ratings along an absolute five-point scale, whereas much of the challenge in projects such as Place Pulse was in reconstructing a ranking from a large set of relative comparisons; this is a basic philosophical difference in approach. The advantage of using absolute scores is that the ratings are always linearly ordered, which is helpful when the expected study sizes are on the smaller side, working with only thousands of responses, or even fewer. We also argue that with high-level and high-complexity concepts such as ‘walkability’ and ‘safety’ the use of relative comparisons is not necessarily helpful, since there are so many aspects of the imagery to consider simultaneously; past psychological research to justify the use of relative comparisons (Stewart et al., 2005) considered only low-complexity concepts such as ‘line lengths’ or ‘sound tones’. Instead of having a complex process of translating relative comparisons into absolute scores, we keep things simple in our system by showing only one image and enabling participants to give a quick first-impression one-click/swipe answer from a selection of five possible responses. This user interface helps participants quickly and smoothly look through many more images than interfaces with a rating slider (Twedt et al., 2016; Larkin et al., 2022) and avoids any biasing problems created by feeding back preliminary machine learning model results into the user interface (Yao et al., 2019), while still allowing participants to able to enjoy rapid progression through the survey.
@@
We were able to easily attract public participants and give them the flexibility to respond at their convenience because our survey is a smartphone-friendly web app rather than a downloaded software application. We did not pay participants (Twedt et al., 2016; Pearson et al., 2024; Larkin et al., 2022; Kruse et al., 2021) but rather relied on the attractiveness of the survey’s appearance as an almost ‘game-like’ app. Our participants took 3.7 sec per image to perform ratings (at the median), which is very close to the 3.6 sec per image timing results reported by Yao et al. (2019), but without the need for feedback (or interference) from a machine learning model.
@@ 5.2 Limitations
Mapillary888mapillary.com/about, as of September 2024 has more than 2 billion images from 190 countries, and especially good coverage in countries like the Netherlands, including almost full panoramic coverage of Amsterdam-area streets. However, although Mapillary has imagery from every continent, the most abundant SVI coverage comes from Europe and North America (Ma et al., 2019). However, commercial SVI is no panacea either: Kim and Jang (2023) analyzed Google SVI coverage of walk commute trajectories in small- to medium-sized cities in the United States and found gaps in nearly half of the routes they investigated; they suggest that researchers consider Mapillary to bolster coverage.
@@
The toolkit is currently split into three separate components, each of which can be used independently of the others. Furthermore, as a web app there are technical reasons for the separation of components: the frontend must integrate with a web server, and the backend must integrate with a database server; both of those servers must be configured correctly as well. However, should a researcher choose to use all three of our components together, this then implies running them all separately, which is less convenient than having a single combined module of some sort. One way around most of this complexity, that we plan to do in the future, is to build a configuration of ‘containers’ to create an application stack (e.g. docker-compose999docs.docker.com/compose/ is popular software for managing multiple containers) with our components plus web server and database server configured within their own containers. However, even that option still requires access to a system installed with said container software and the necessary networking set-up and capabilities to serve web content on the Internet. If a researcher or citizen scientist does not have access to such a system, then it is possible to purchase access to one from a selection of many application server hosting providers.
@@ 6 Conclusion
We offer to the research community a free and open-source toolkit for downloading, processing and filtering VSVI from a given geographic region, and deploying a mobile-friendly human perception survey web app with a specially customized user interface on the resulting images. In the spirit of citizen science and the FAIR research principles, anyone may easily clone, modify and deploy this toolkit on any location of interest. We integrated it with the Mapillary platform because they provide open access to VSVI from many cities around the world, including some with little or no Google Street View coverage as of this writing such as Casablanca, Lusaka and Managua. We anticipate that this toolkit will be used to build deep learning models with the collected data and make predictions about human perception of large amounts of SVI over a wider area. Ultimately, we hope to see such perception data used in research that can help guide future planning and development choices, and therefore improve the quality of life for many people in urban areas.
@@ Appendix C Processing and filtering VSVI details
For full up-to-date usage information please see the Spatial Data Science and Geo AI Lab101010github.com/Spatial-Data-Science-and-GEO-AI-Lab/percept web site.
@@ C.1 Downloading imagery from Mapillary
This script takes a Mapillary API key111111www.mapillary.com/developer and a bounding box (west, south, east, north) and conducts a lengthy but robust and restartable procedure to methodologically find and download each tile data file, cache it, and then download all of the eligible SVI that is found within the tile (and within the bounding box).